# Category-free Out-of-Distribution Node Detection with Feature Resonance

### v1

@@ Appendix F Experiment Details
Dataset Squirrel WikiCS YelpChi Amazon Reddit
Learning rate ($\alpha$) 0.005 0.01 0.005 0.005 0.01
$h_{\theta}$ layers 1 1 1 1 1
$g_{\theta}(\cdot)$ layers 2 2 2 2 2
Hidden states 16 16 16 16 16
Dropout rate 0.1 0.1 0.1 0.1 0.1
$\mathrm{n}$ 2 1 2 2 1
$\lambda$ 0.5 0.5 0.5 0.5 0.5
F.1 Hyperparameter
@@ F.3 Dataset Description
To thoroughly evaluate the effectiveness of RSL, we perform experiments on five diverse and real-world OOD node detection datasets:
@@
  • Squirrel (Rozemberczki et al., 2021): A Wikipedia network where nodes correspond to English Wikipedia articles, and edges represent mutual hyperlinks. Nodes are categorized into five classes following Geom-GCN (Pei et al., 2020) annotations, with the network exhibiting a high level of heterophily.
@@
  • WikiCS (Mernyei & Cangea, 2020): This dataset consists of nodes representing articles in the Computer Science domain. Edges are based on hyperlinks, and nodes are classified into 10 categories, each corresponding to a unique sub-field of Computer Science.
@@
  • YelpChi (Rayana & Akoglu, 2015): Derived from Yelp, this dataset includes hotel and restaurant reviews. Legitimate reviews are labeled as ID nodes, while spam reviews are considered OOD nodes.
@@
  • Amazon (McAuley & Leskovec, 2013): Contains reviews from the Musical Instrument category on Amazon.com. ID nodes represent benign users, while OOD nodes correspond to fraudulent users.
@@
  • Reddit (Kumar et al., 2019): A dataset comprising user posts collected from various subreddits over a month. Normal users are treated as ID nodes, while banned users are labeled as OOD nodes.
@@
We follow the same data preprocessing steps as EnergyDef (Gong & Sun, 2024). Both Squirrel and WikiCS datasets are loaded using the DGL (Wang et al., 2019) package. For Squirrel, class {1} is selected as the OOD class, while {0, 2, 3, 4} are designated as ID classes. In the case of WikiCS, {4, 5} are chosen as OOD classes, with the remaining eight classes treated as ID. The YelpChi and Amazon datasets are processed based on the methodology described in (Dou et al., 2020), and the Reddit dataset is prepared using the PyGod (Liu et al., 2022) package.
F.4 Baseline Description
@@
  • LOF-KNN (Breunig et al., 2000) calculates the OOD scores of node attributes by assessing the deviation in local density relative to the k-nearest node attributes.
@@
  • GCNAE (Kipf & Welling, 2016b) swaps the MLP backbone for a GCN in the autoencoder. The OOD score is determined in the same way as MLPAE, following the same training process.
@@
  • GAAN (Chen et al., 2020) is a generative adversarial network for attributes that evaluates sample reconstruction error and the confidence of recognizing real samples to predict OOD nodes.
@@
  • DOMINANT (Ding et al., 2019) combines a structure reconstruction decoder and an attribute reconstruction decoder. The total reconstruction error for each node consists of the errors from both decoders.
@@
  • ANOMALOUS (Peng et al., 2018) is an anomaly detection method that utilizes CUR decomposition and residual analysis for identifying OOD nodes.
@@
  • SL-GAD (Zheng et al., 2021) derives OOD scores for nodes by considering two aspects: reconstruction error and contrastive scores.
@@
  • GOAD (Bergman & Hoshen, 2020) enhances training data by transforming it into independent spaces and trains a classifier to align the augmented data with the corresponding transformations. OOD scores are then calculated based on the distances between OOD inputs and the centers of the transformation spaces. For graph-structured data, we use the same GNN backbone as EnergyDef-h.
@@
  • NeuTral AD (Qiu et al., 2021) uses learnable transformations to embed data into a semantic space. The OOD score is determined by a contrastive loss applied to the transformed data.
@@
  • GKDE (Zhao et al., 2020) predicts Dirichlet distributions for nodes and derives uncertainty as OOD scores by aggregating information from multiple sources.
@@
  • OODGAT (Song & Wang, 2022) is an entropy-based OOD detector that assumes node category labels are available. It uses a Graph Attention Network as the backbone and determines OOD nodes based on category distribution outcomes.
@@
  • GNNSafe (Wu et al., 2023) calculates OOD scores by applying the LogSumExp function over the output logits of a GNN classifier, which is trained with multi-category labels. The rationale for the OOD score is the similarity between the Softmax function and the Boltzmann distribution.
@@
  • SSD (Sehwag et al., 2021) is an outlier detector that leverages self-supervised representation learning and Mahalanobis distance-based detection on unlabeled ID data. We use twice dropout to generate positive pairs for contrastive learning like SimCSE (Gao et al., 2021).
@@
  • EnergyDef (Gong & Sun, 2024) uses Langevin dynamics to generate synthetic OOD nodes for training the OOD node classifier.
F.5 Implementation Details
@@
We adopt the same dataset settings as EnergyDef (Gong & Sun, 2024). It is worth noting that, under this dataset setup, the features of unknown nodes are accessible. Therefore, using the features of unknown nodes during the training phase to filter reliable OOD nodes is a legitimate strategy. Specifically, for the Squirrel and WikiCS datasets, we randomly select one and two classes as OOD classes, respectively. In the case of fraud detection datasets, we categorize a large number of legitimate entities as ID nodes and fraudsters as OOD nodes. We allocate 40% of the ID class nodes for training, with the remaining nodes split into a 1:2 ratio for validation and testing, ensuring stratified random sampling based on ID/OOD labels.
@@
We report the average value of five independent runs for each dataset. The hyper-parameters are shown in Table 8. The anomaly detection baselines are trained entirely based on graph structures and node attributes without requiring ID annotations. We adapt these models to the specifications of our OOD node detection tasks by minimizing the corresponding loss items solely on the ID nodes, where applicable.
Appendix G More Experiments
(a) Reddit
(b) Squirrel
(c) YelpChi
(d) WikiCS
G.1 The Feature Resonance Phenomenon Induced by Different Target Vectors
@@
We explore the phenomenon of feature resonance using different target vectors. Experiments are conducted on two datasets with real $N$-category labels, Squirrel and WikiCS ($N$ represents the number of categories). First, based on the neural collapse theory (Papyan et al., 2020; Zhou et al., 2022), we preset $N$ target vectors, each representing a category. These $N$ target vectors form an equiangular tight frame, maximizing the separation between them. The definition of the simplex equiangular tight frame is introduced as follows:
Definition G.1.
@@
Simplex ETF. (Xiao et al., 2024) A simplex equiangular tight frame (ETF) refers to a collection of K equal-length and maximally-equiangular P-dimensional embedding vectors $\mathbf{E}=[e_{1},\cdots,e_{K}]\in\mathbb{R}^{P\times K}$ which satisfies:
@@
where $\mathbf{I}_{K}$ is the identity matrix,$\mathbf{1}_{K}$ is an all-ones vector, and $\mathbf{U}\in\mathbb{R}^{P\times K}(P\geq K)$ allows a rotation.
All vectors in a simplex ETF $\mathbf{E}$ have an equal $\mathit{l}_{2}$ norm and the same pair-wise maximal equiangular angle $-\frac{1}{K-1}$,
@@
where $\delta_{k_{1},k_{2}}=1$ when $k_{1}=k_{2}$ and 00 otherwise.
@@
$\ell(h_{\theta_{t}}({\boldsymbol{X}}_{\text{known}}),e)=\mathbb{E}(\parallel\mathbf{E}_{\text{known}}-({\boldsymbol{X}}_{\text{known}}\mathbf{W}^{\top})\parallel^{2}_{2})$ (21)
@@
The trajectory trends and lengths of unknown ID nodes differ significantly from those of OOD nodes, with the former showing more distinct trends and longer trajectories. We refer to this as the feature resonance phenomenon and leverage it to filter OOD nodes. As shown in Table 4, under the “True multi-label” row, the experimental results demonstrate that this method is effective and performs well. Interestingly, even with random labels for known ID nodes or aligning all known ID representations to a fixed target vector, unknown ID nodes consistently exhibit longer trajectories than unknown OOD nodes, as shown in Table 4.
@@
(a) WikiCS
(b) YelpChi
@@
Figure 7: The impact of different sliding window widths on the performance of detecting OOD nodes. When the width is 1, it corresponds to the resonance-based score $\tau$.
(a) Pre-training (EnergyDef)
(b) Pre-training (Ours)
(c) Post-training (EnergyDef)
(d) Post-training (Ours)
@@
Figure 8: T-SNE visualization of node embeddings on the dataset WikiCS. (a) Synthetic nodes (red) generated by EnergyDef fail to accurately represent the actual features of OOD nodes (blue), whereas ours can, as shown in (b). (c) Representations of ID (green) and OOD (blue) nodes trained with synthetic nodes generated by EnergyDef are poorly separated, whereas ours can, as shown in (d).
G.2 Variation of Microscopic Feature Resonance During Training

### v2

@@ Appendix E Experiment Details
Dataset Squirrel WikiCS YelpChi Amazon Reddit Cora Citeseer Pubmed Chameleon
Learning rate ($\alpha$) 0.005 0.01 0.005 0.005 0.01 0.005 0.005 0.01 0.005
$h_{\theta}$ layers 1 1 1 1 1 1 1 1 1
$g_{\theta}(\cdot)$ layers 2 2 2 2 2 2 2 2 2
Hidden states 16 16 16 16 16 16 16 16 16
Dropout rate 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1
$\mathrm{n}$ 2 1 2 2 1 10 10 5 2
$\lambda$ 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5
E.1 Hyperparameter
@@ E.3 Dataset Description
To thoroughly evaluate the effectiveness of RSL, we perform experiments on real-world node-level and graph-level OOD detection datasets:
  • Node-level Datasets:
@@
    • Squirrel [Rozemberczki et al., 2021]: A Wikipedia network where nodes correspond to English Wikipedia articles, and edges represent mutual hyperlinks. Nodes are categorized into five classes following Geom-GCN [Pei et al., 2020] annotations, with the network exhibiting a high level of heterophily.
@@
    • WikiCS [Mernyei and Cangea, 2020]: This dataset consists of nodes representing articles in the Computer Science domain. Edges are based on hyperlinks, and nodes are classified into 10 categories, each corresponding to a unique sub-field of Computer Science.
@@
    • YelpChi [Rayana and Akoglu, 2015]: Derived from Yelp, this dataset includes hotel and restaurant reviews. Legitimate reviews are labeled as ID nodes, while spam reviews are considered OOD nodes.
@@
    • Amazon [McAuley and Leskovec, 2013]: Contains reviews from the Musical Instrument category on Amazon.com. ID nodes represent benign users, while OOD nodes correspond to fraudulent users.
@@
    • Reddit [Kumar et al., 2019]: A dataset comprising user posts collected from various subreddits over a month. Normal users are treated as ID nodes, while banned users are labeled as OOD nodes.
    • Chameleon Rozemberczki et al. [2021] is a Wikipedia network with 5 classes, where nodes represent web pages and edges represent hyperlinks between them. Node features represent several informative nouns in the Wikipedia pages, and the task is to predict the average daily traffic of the web page Fey and Lenssen [2019].
    • Cora [Kipf and Welling, 2016a] is a citation graph with 2,708 nodes, 5,429 edges, 1,433 features, and 7 classes, widely used for node classification and link prediction. Under the Label Leave-out setting, 3 classes are treated as ID and 4 as OOD.
    • Citeseer [Kipf and Welling, 2016a] contains 3,327 nodes, 4,732 edges, 3,703 features, and 6 classes. We apply the same OOD generation strategies as above, designating 3 classes as ID and 3 as OOD under the Label Leave-out setting.
    • PubMed [Kipf and Welling, 2016a], a biomedical citation graph, includes 19,717 nodes, 44,338 edges, 500 features, and 3 classes. We follow the same OOD generation and semi-supervised training procedure, using 2 classes as ID and 1 as OOD under the Label Leave-out setting.
@@
    For the Squirrel, WikiCS, YelpChi, Amazon, and Reddit datasets, we follow the same data preprocessing steps as EnergyDef [Gong and Sun, 2024]. Both Squirrel and WikiCS datasets are loaded using the DGL [Wang et al., 2019] package. For Squirrel, class {1} is selected as the OOD class, while {0, 2, 3, 4} are designated as ID classes. In the case of WikiCS, {4, 5} are chosen as OOD classes, with the remaining eight classes treated as ID. The YelpChi and Amazon datasets are processed based on the methodology described in [Dou et al., 2020], and the Reddit dataset is prepared using the PyGod [Liu et al., 2022] package. For the Cora and Chameleon datasets, we follow the data processing procedure used in GRASP [Ma et al., 2024].
  • Graph-level Datasets:
    • ENZYMES Morris et al. [2020] is a graph dataset constructed based on the structural properties of protein molecules. It contains a total of 600 graphs, each representing one protein sample, across six different classes. The dataset includes 19,580 nodes and 174,564 edges, with each node having a feature vector of dimension 3.
    • PROTEINS Morris et al. [2020] is a dataset of proteins that are classified as enzymes or non-enzymes. The dataset includes 1,113 graphs.
    • ClinTox [Wu et al., 2018] compares drugs approved by the FDA and drugs that have failed clinical trials for toxicity reasons. The dataset includes two classification tasks for 1491 drug compounds with known chemical structures: (1) clinical trial toxicity (or absence of toxicity) and (2) FDA approval status.
    • LIPO [Wu et al., 2018] is a dataset included in MoleculeNet [Wu et al., 2018]. It measures the experimental results of octanol/water distribution coefficient(logD at pH 7.4).
    We follow the data processing procedure used in GOOD-D [Liu et al., 2023] that 90% of ID samples are used for training, and 10% of ID samples and the same number of OOD samples are integrated together for testing.
E.4 Baseline Description
  • Node-level Baselines:
@@ E.4 Baseline Description
    • LOF-KNN [Breunig et al., 2000] calculates the OOD scores of node attributes by assessing the deviation in local density relative to the k-nearest node attributes.
@@
    • GCNAE [Kipf and Welling, 2016b] swaps the MLP backbone for a GCN in the autoencoder. The OOD score is determined in the same way as MLPAE, following the same training process.
@@
    • GAAN [Chen et al., 2020] is a generative adversarial network for attributes that evaluates sample reconstruction error and the confidence of recognizing real samples to predict OOD nodes.
@@
    • DOMINANT [Ding et al., 2019] combines a structure reconstruction decoder and an attribute reconstruction decoder. The total reconstruction error for each node consists of the errors from both decoders.
@@
    • ANOMALOUS [Peng et al., 2018] is an anomaly detection method that utilizes CUR decomposition and residual analysis for identifying OOD nodes.
@@
    • SL-GAD [Zheng et al., 2021] derives OOD scores for nodes by considering two aspects: reconstruction error and contrastive scores.
@@
    • GOAD [Bergman and Hoshen, 2020] enhances training data by transforming it into independent spaces and trains a classifier to align the augmented data with the corresponding transformations. OOD scores are then calculated based on the distances between OOD inputs and the centers of the transformation spaces. For graph-structured data, we use the same GNN backbone as EnergyDef-h.
@@
    • NeuTral AD [Qiu et al., 2021] uses learnable transformations to embed data into a semantic space. The OOD score is determined by a contrastive loss applied to the transformed data.
    • MSP Hendrycks and Gimpel [2016]: Uses the maximum softmax probability as the OOD score. The method is simple but has limited performance on models with high confidence.
    • ODIN Liang et al. [2017]: Improves OOD detection by temperature scaling and input perturbation, but is sensitive to hyperparameters.
    • Mahalanobis Lee et al. [2018]: Calculates the feature distance between a sample and ID data based on Mahalanobis distance, suitable for scenarios assuming a Gaussian distribution.
    • OE Hendrycks et al. [2018]: Optimizes using additional OOD data during training, relying on the availability of OOD data.
    • Energy & Energy FineTune Liu et al. [2020]: Uses an energy function instead of softmax probabilities for OOD scoring, and can improve detection performance by fine-tuning with OOD data.
@@
    • GKDE [Zhao et al., 2020] predicts Dirichlet distributions for nodes and derives uncertainty as OOD scores by aggregating information from multiple sources.
    • GPN Stadler et al. [2021]: Based on Bayesian posterior inference, performs OOD detection through uncertainty estimation. It is suitable for graph data but sensitive to hyperparameters.
@@
    • OODGAT [Song and Wang, 2022] is an entropy-based OOD detector that assumes node category labels are available. It uses a Graph Attention Network as the backbone and determines OOD nodes based on category distribution outcomes.
@@
    • GNNSafe [Wu et al., 2023] calculates OOD scores by applying the LogSumExp function over the output logits of a GNN classifier, which is trained with multi-category labels. The rationale for the OOD score is the similarity between the Softmax function and the Boltzmann distribution.
    • NodeSafe [Yang et al., 2025] reduces the occurrence of extreme energy values by enforcing consistency in the logit norms, thereby decreasing the variance within both the ID and OOD energy distributions, which enhances the performance of OOD node detection.
    • GRASP [Ma et al., 2024] enhances OOD node detection performance by amplifying the graph’s homophily through rewiring, thereby improving the effect of score propagation.
@@
    • SSD [Sehwag et al., 2021] is an outlier detector that leverages self-supervised representation learning and Mahalanobis distance-based detection on unlabeled ID data. We use twice dropout to generate positive pairs for contrastive learning like SimCSE [Gao et al., 2021].
@@
    • EnergyDef Gong and Sun [2024] uses Langevin dynamics to generate synthetic OOD nodes for training the OOD node classifier.
  • Graph-level Baselines:
    • Graph kernel + detector: This category of methods involves two main steps: first, graph kernel techniques are employed to transform graphs into vector-based features Vishwanathan et al. [2010]; next, out-of-distribution (OOD) detection algorithms are applied to these feature vectors. Specifically, we adopt the Weisfeiler-Lehman (WL) kernel Shervashidze et al. [2011] and the propagation kernel (PK) Neumann et al. [2016] for representation, combined with anomaly detectors such as local outlier factor (LOF) Breunig et al. [2000], one-class SVM (OCSVM) Manevitz and Yousef [2001], and isolation forest (iF) Liu et al. [2008].
    • GCL + detector: These approaches leverage recent advances in Graph Contrastive Learning (GCL) to derive graph-level embeddings, which are then assessed by OOD detection methods. We employ two representative GCL techniques—InfoGraph Sun et al. [2020] and GraphCL You et al. [2020]—to generate node or graph representations. For detecting OOD instances, we consider both the isolation forest (iF) Liu et al. [2008] and a Mahalanobis distance-based (MD) detector, which has demonstrated strong performance in identifying OOD data Sehwag et al. [2021], Zhou et al. [2021]. GOOD-D Liu et al. [2023] can capture the latent ID patterns and accurately detect OOD graphs based on the semantic inconsistency in different granularities by performing hierarchical contrastive learning on the augmented graphs.
    • End-to-end: We also evaluate our model against end-to-end graph anomaly detection baselines. One such approach is OCGIN Zhao and Akoglu [2023], which utilizes a GIN encoder trained with a support vector data description (SVDD) loss. Another is GLocalKD Ma et al. [2022], which detects anomalous samples through a knowledge distillation framework.
E.5 Implementation Details
@@
We adopt the same dataset settings as EnergyDef [Gong and Sun, 2024], and we use GCN [Kipf and Welling, 2016a] as the encoder. It is worth noting that, under this dataset setup, the features of unknown nodes are accessible. Therefore, using the features of unknown nodes during the training phase to filter reliable OOD nodes is a legitimate strategy. Specifically, for the Squirrel and WikiCS datasets, we randomly select one and two classes as OOD classes, respectively. In the case of fraud detection datasets, we categorize a large number of legitimate entities as ID nodes and fraudsters as OOD nodes. We allocate 40% of the ID class nodes for training, with the remaining nodes split into a 1:2 ratio for validation and testing, ensuring stratified random sampling based on ID/OOD labels.
@@
We report the average value of five runs for each dataset. The hyper-parameters are shown in Table 10. The anomaly detection baselines are trained entirely based on graph structures and node attributes without requiring ID annotations. We adapt these models to the specifications of our OOD node detection tasks by minimizing the corresponding loss items solely on the ID nodes, where applicable.
Appendix F More Experiments
(a) The performance of using resonance-based score $\tau$ to detect OOD nodes varies with training progress. The higher the AUROC, the better, and the lower the FPR95, the better.
F.1 The Feature Resonance Phenomenon Induced by Different Target Vectors
@@
We explore the phenomenon of feature resonance using different target vectors. Experiments are conducted on two datasets with real $N$-category labels, Squirrel and WikiCS ($N$ represents the number of categories). First, based on the neural collapse theory [Papyan et al., 2020, Zhou et al., 2022], we preset $N$ target vectors, each representing a category. These $N$ target vectors form an equiangular tight frame, maximizing the separation between them. The definition of the simplex equiangular tight frame is introduced as follows:
Definition 6.
@@
Simplex ETF. [Xiao et al., 2024] A simplex equiangular tight frame (ETF) refers to a collection of K equal-length and maximally-equiangular P-dimensional embedding vectors $\mathbf{E}=[e_{1},\cdots,e_{K}]\in\mathbb{R}^{P\times K}$ which satisfies:
@@
[…]
All vectors in a simplex ETF $\mathbf{E}$ have an equal $\mathit{l}_{2}$ norm and the same pair-wise maximal equiangular angle $\frac{1}{K-1}$,
@@
where $\delta_{k_{1},k_{2}}=1$ when $k_{1}=k_{2}$ and 0 otherwise.
@@
$\ell(h_{\theta_{t}}({\bm{X}}_{\text{known}}),e)=\mathbb{E}(\parallel\mathbf{E}_{\text{known}}-({\bm{X}}_{\text{known}}\mathbf{W}^{\top})\parallel^{2}_{2})$ (18)
@@
The trajectory trends and lengths of unknown ID nodes differ significantly from those of OOD nodes, with the former showing more distinct trends and longer trajectories. We refer to this as the feature resonance phenomenon and leverage it to filter OOD nodes. As shown in Table 6, under the “True multi-label" row, the experimental results demonstrate that this method is effective and performs well. Interestingly, even with random labels for known ID nodes or aligning all known ID representations to a fixed target vector, unknown ID nodes consistently exhibit longer trajectories than unknown OOD nodes, as shown in Table 6.
@@
Figure 7: Performance of detecting OOD nodes with different metrics. $\tau$ represents the resonance-based score, the “Overall Trajectory" represents the total cumulative length of the training trajectory $\hat{F}(\tilde{\mathbf{x}}_{i})=\sum_{t}\tau_{i}$, and the “Sliding Window" refers to the cumulative $\tau$ within a window of width 10: $\hat{F}_{10}(\tilde{\mathbf{x}}_{i})=\sum^{t}_{t-10}\tau_{i}$.
@@
(a) The impact of different sliding window widths on the performance of detecting OOD nodes. When the width is 1, it corresponds to the resonance-based score $\tau$.
(b) The score distribution of ID nodes and OOD nodes on Amazon.
@@
(c) T-SNE visualization of node embeddings on the dataset WikiCS. (a) Synthetic nodes (red) generated by EnergyDef fail to accurately represent the actual features of OOD nodes (blue), whereas ours can, as shown in (b). (c) Representations of ID (green) and OOD (blue) nodes trained with synthetic nodes generated by EnergyDef are poorly separated, whereas ours can, as shown in (d).
F.2 Variation of Microscopic Feature Resonance During Training